OpenStudy (anonymous):

A 1180 kg car is pulling a 165 kg trailer along a level road. Friction is negligiable. The car accelerates with an acceleration of 1.30 m/s2. What is the force exerted by the car on the trailer?

OpenStudy (anonymous):

165kg*1.3m/s^2 = 214.5N

OpenStudy (sriram):

force = mass * acc. mass of trailer=165 kg its acc=acc of car=1.3m/s2 hence force= 165*1.3 =214.5 N